# Language Models Trained to do Arithmetic Predict Human Risky and Intertemporal Choice

### v1

@@ 5 Experimental Results
5.1 Model Comparisons
In this section, we present the experimental results from our model comparisons. We first obtained embeddings from the Arithmetic-GPT model, evaluating versions of the model that were pretrained on each of our four distinct synthetic datasets as well as a version without any training. Specifically, we extracted embeddings for the expected values of the two options, denoted as $e_{A}$ and $e_{B}$. Additionally, we obtained embeddings for the difference in expected values, denoted as $e_{A-B}$. All embeddings were derived from the representation in the final layer before the autoregressive prediction. We then performed a logistic regression using $e_{A}$, $e_{B}$, and $e_{A-B}$ as independent variables, with human choice probabilities as the dependent variable. The adjusted $R^{2}$ results of the logistic regressions are summarized in the first 5 rows of Table 3. These results indicate that the embeddings from the Arithmetic-GPT model pretrained on ecologically valid synthetic datasets most accurately capture human choices. This model also outperforms the embeddings obtained from the LLaMA-3-70b-Instruct model (the 7th row of Table 3), suggesting that pretraining on synthetic datasets is sufficient to create a strong correspondence between LLMs and human decision-making. The log probabilities from the same LLaMA3 model perform poorly in comparison to human data (the 6th row of Table 3), replicating previous findings using choice rates reported from LLaMA1 [3].
To benchmark the performance of Arithmetic-GPT models in explaining human data, we directly fitted behavioral models and MLPs on each of the human choice datasets. The CPT and hyperbolic discounting models are leading behavioral models in risky and intertemporal choices, respectively [21, 22]. The behavioral models have interpretable mechanisms and contain few free parameters. However, they do not explain human data as well as the embeddings from both LLaMA3 and Arithmetic-GPT. In contrast, fitting an MLP directly on human data potentially reveals the ceiling performance of any model in explaining the data [c.f. 2]. Except for the intertemporal choice tasks, MLPs outperform all other models. It is important to note that MLP training was based on task features rather than text descriptions that simulated participants’ experiences or arithmetic equations that mimic a rational agent’s computations. Consequently, these differences in input formats could also lead to diverging performance. The arithmetic format of intertemporal choice may provide a better fit for human data. Indeed, there is increasing evidence from experimental economics suggesting that the complexity of discounting values, even in the absence of actual payoff delays, influences intertemporal choices [10]. We include a robustness check using 10-fold cross-validation in Appendix B.
Table 3: Variances of human choices ($R^{2}$) explained by the embeddings from the pretrained Arithmetic-GPT model compared to other computational models.
@@ 5.1 Model Comparisons
Model Training data choice13k cpc18 gershman20 agrawal23
@@
LLaMA3(arith.)h Undisclosedg 63.6% 34.8% 69.3% 96.0%
CPTi Human choices 51.5% 54.2% N/A N/A
@@
Note. aUniform synthetic datasets. bThe same uniform synthetic datasets but with the answers on the right-hand side of the equations removed and the signs randomized between positive and negative. cEcological synthetic datasets. dThe same ecological synthetic datasets but with the answers on the right-hand side of the equations removed and the signs randomized between positive and negative. eEmbeddings from an untrained Arithmetic-GPT model. fLog probabilities of LLaMA3 elicited from text descriptions of choice problems. gThe training data is not publicly available, but Meta has disclosed some summary statistics of the training corpus. hEmbeddings from LLaMA3 elicited from arithmetic equations of choice problems. iCumulative prospect theory for risky choices [21]. jThe hyperbolic discounting model for intertemporal choices [22].
5.2 Implicit functions of probabilities, values, and discount factors from Arithmetic-GPT
@@
We find that the embeddings from ecologically pretrained Arithmetic-GPT replicate classical findings from behavioral economics, including value and probability weighting functions from the cumulative prospect theory [40] and the hyperbolic discounting function [22]. Specifically, probabilities close to 0.5 are more similar to each other than to probabilities close to either 0 or 1 (see Figure 2A and Equation 4). Embeddings for values illustrate concavity for positive values, convexity for negative values, and a steeper slope for negative values than for positive values (see Figure 2B and Equation 3). These features reflect risk aversion for gains, risk seeking for losses, and loss aversion [21, 40]. Moreover, embeddings for the discount factor demonstrate that distant times are more similar (see Figure 2C and Equation 6), enabling a present-bias [29].
Figure 2: Embeddings from ecologically pretrained Arithmetic-GPT for inputs including (A) probabilities, (B) values, and (C) discount factors. Inputs are shown along the horizontal axes and embeddings are shown on the vertical axes. The embeddings, shown as black dots, were reduced to 1D using multidimensional scaling. Embeddings for probabilities and discount factors are normalized between 0 and 1. The red curves represent the best-fitting behavioral economic models: (A) the probability weighting function from CPT with best-fitting $\gamma=0.58$, (B) the utility function from CPT with best-fitting $\alpha=0.42,\beta=0.45,\lambda=1.4$, and (C) the hyperbolic discount function with best-fitting $k=0.08$.
For the probability weighting function, the curvature parameter ($\gamma$) of human participants was found to be 0.61 for gains and 0.69 for losses [40]. Follow-up replications by [44] found $\gamma$ values around 0.7. Moreover, the utility curvature parameters ($\alpha$ and $\beta$) typically range between 0.5 and 0.9, while the loss aversion parameter ($\lambda$) is approximately 2.25 [40]. Regarding human intertemporal choices, typical $k$ values range between 0.01 and 0.1 for small magnitude studies [32]. Comparing to these best-fitting parameters from human experiments, we observe that the implicit functions derived from the embeddings of Arithmetic-GPT also quantitatively match those observed in humans (see Figure 2). The implicit functions, however, exhibit discontinuities, suggesting that the smooth functions derived from human theories may not be directly applicable to explain the embeddings of Arithmetic-GPT.
@@ 6 Discussion
We introduced an approach to transforming language models into cognitive models by pretraining them on tasks that mirror the computational process performed by a rational agent. Specifically, we pretrained Arithmetic-GPT to calculate expected values using ecologically distributed probabilities and values. This pretraining allows the model to better capture human risky and intertemporal choices compared to classical behavioral models and, in some cases, even surpass the performance of the general-purpose LLaMA3 model. These results have implications for a number of questions in cognitive science and machine learning, although we also note the limitations of our current work.
@@
Language Models as Cognitive Models. There is a growing research effort focused on exploring LLMs as scientific tools for understanding the human mind [3, 13, 18]. Despite the challenges outlined in Section 2, LLMs enable researchers to investigate cognitive processes in ways that are not feasible with human participants. Recent studies have even demonstrated that fine-tuning a LLaMA1 model on human data allows the model to outperform classical models in explaining this data [3]. However, while the LLaMA series’ model architecture and weights are publicly available, researchers lack access to the training data, which makes crucial scientific inference practically impossible. In contrast, we explicitly manipulate the training data for our Arithmetic-GPT, thereby uncovering a key factor that contributes to the model’s ability to explain human choices.
@@
Bayesian Models of Cognition, Meta-learning, and Pre-training. Bayesian models of cognition have been instrumental in understanding human performance across a variety of cognitive tasks by providing optimal solutions to the inductive inference problem. Recently, it has been argued that Bayesian models and neural network models can be viewed as complementary to one another [17]. Neural networks that are meta-trained have also been shown to exhibit properties consistent with Bayesian models [23, 28]. Moreover, pretraining a neural network model for improved performance on downstream tasks can be seen as a form of meta-learning or the acquisition of useful inductive biases, similar to Bayesian models [19]. Our work makes the implicit priors learned by Arithmetic-GPT more explicit by specifying the synthetic dataset on which it was pre-trained.
@@
Computationally Equivalent Tasks for Cognitive Modeling. Modeling human cognition is challenging because the hypothesis space of possible cognitive mechanisms that can explain human data equally well is vast. This makes principles of rationality desirable, as assuming people are rational in some sense greatly constrains the hypothesis space, thereby making scientific inference more effective. However, human rationality has been a subject of debate in economics and psychology for over a century [42, 16, 21, 25, 37]. While significant progress has been made in understanding human rationality [e.g., 25, 15], the advent of LLMs seems to challenge the need for rational theories. Simply training LLMs to predict the next word appears sufficient to produce human-like behaviors, suggesting that we can model human cognition without the constraints imposed by rational theories. However, our experimental results suggest an alternative route to enhancing the correspondence between behaviors produced by LLMs and humans: pretraining on computationally equivalent tasks that a rational agent would need to master. Future research should investigate the impact of different assumptions about the nature of rationality on task content and distributions, and explore whether there are more effective assumptions for pre-training models to explain human behavior.
@@
Implications for Theories of Human Risk and Time Preferences. The success of Arithmetic-GPT in explaining human risky and intertemporal choices has significant implications for theoretical work on human risk and time preferences. While the Homo economicus portrayal of human beings as perfectly rational and self-interested agents has been inadequate in describing human choices [16, 21], existing behavioral models that deviate from rationality do not generalize well across different task domains. For instance, it is challenging to use CPT to model time preferences or to use the hyperbolic discounting model to explain risk preferences. In contrast, Arithmetic-GPT has demonstrated substantial transferability across task domains. The same model, in principle, can be adapted to other judgment and decision-making tasks such as social choice, probability judgment, and belief updating. The key factor enabling language models to generalize across a wide range of tasks is the presence of a language interface, which underpins a significant range of cognitive tasks.
@@
The Importance of Training Data Disclosure. Our results demonstrate that the training data used for LLMs is crucial for understanding their emergent capabilities and the inductive biases they acquire during pretraining. Adjusting the distribution of, or ablating, the training data can significantly affect the degree to which LLMs correspond with human behaviors. These findings suggest that existing off-the-shelf LLMs, whether proprietary or open-weight, including the GPT series [6] and the LLaMA series [39], are less effective as models of human cognition. This is primarily because their training data is rarely disclosed, making it difficult for scientists to precisely identify the sources of human-like behaviors in these models.
@@
Limitations and Future Research. While we have made progress in addressing many challenges associated with using LLMs as cognitive models, some issues remain unresolved. To address the data gap between LLMs and human learners, we limited the scope of our synthetic dataset to the arithmetic of expected value calculations. Despite this, LLMs still require a substantial amount of training data to perform arithmetic accurately within a limited range of input values. Further research is needed to continue bridging this data gap.
@@
Many additional ablation studies could be performed on model architectures and training objectives. Our work is fundamentally limited to autoregressive training and a decoder-only transformer architecture. We believe that alternative training mechanisms and model architectures could potentially yield better embeddings from LLMs. One area for future research is estimating the lower bounds on model size necessary to achieve a certain level of correspondence with human behavior.
@@ Appendix A Prompts
System ‘‘You are participating in a gambling game where you will be shown two options, Gamble A and Gamble B. Your task is to choose between the two. You must select one option.’’
User ‘‘Gamble A offers a 10% chance to win \$10 and a 90% chance to lose \$12. Gamble B offers a 40% chance to lose \$13 and a 60% chance to win \$22. Please limit your answer to either ‘A’ or ‘B’.’’
@@
System ‘‘You are participating in an intertemporal choice experiment. You will be presented with two options, A and B, and your task is to choose between them. You must select one option.’’
User ‘‘Option A offers \$80 after 30 days. Option B offers \$13 immediately. Please limit your answer to either ‘A’ or ‘B’.’’
@@ Appendix B Robustness Check
We supplement our main results with a robustness check using 10-fold cross-validation. Each dataset of human choices was randomly partitioned into a 90% training set and a 10% validation set. For the gershman20 dataset, a stratified train/validation split was employed due to the certainty equivalence design of the experiment [14], ensuring that choices made within the same problem were grouped together.
Embeddings from Arithmetic-GPT models and LLaMA-3-70b-Instruct were fitted using logistic regression with a LASSO penalty on the training set. MLPs were fitted to task features in the training set. Performance for all models was assessed on the validation set, with $R^{2}$ calculated as the squared Pearson’s $r$. This process was repeated 10 times, each with a random train/validation split. The mean and standard error (SE) of $R^{2}$ are summarized in Table 4, demonstrating a replication of our main results presented in Table 3.
Table 4: Cross-validation results. Variances of human choices ($R^{2}$) on the validation set explained by models. Numbers in parentheses represent standard errors from 10-fold cross-validation.
@@
Model Training data choice13k cpc18 gershman20 agrawal23
Arith.-GPT Synthetic (unif.) 55.9% (1.4%) 46.3% (3.1%) 56.7% (1.1%) 81.5% (0.4%)
@@
Appendix C Implementation Details

### v2

@@ 5 Experimental Results
Table 3: Proportion of the variance of human choices explained ($R^{2}$) by embeddings from the pretrained Arithmetic-GPT model compared to other computational models. Bold numbers indicate the best models within each group.
@@
Model Training data choices13k cpc18 gershman20 agrawal23
@@
LLaMA3 (txt emb.)h Undisclosedg 66.2% 55.7% 66.5% 98.0%
LLaMA3 (arith.)i Undisclosedg 63.6% 34.8% 69.3% 96.0%
Prospect theory Human choices 51.5% 54.2% N/A N/A
@@
Note. aUniform synthetic datasets. bThe same uniform synthetic datasets but with the answers on the right-hand side of the equations removed and the signs randomized between positive and negative. cEcological synthetic datasets. dThe same ecological synthetic datasets but with the answers on the right-hand side of the equations removed and the signs randomized between positive and negative. eEmbeddings from an untrained Arithmetic-GPT model with randomly initialized weights. fLog probabilities of LLaMA3 elicited from text descriptions of choice problems. gThe training data is not publicly available, but Meta has disclosed some summary statistics of the training corpus. hEmbeddings of LLaMA3 elicited from text descriptions of choice problems. iEmbeddings of LLaMA3 elicited from arithmetic equations of choice problems. jThe hyperbolic discounting model for intertemporal choices. kLogistic regression results when applied to the expected value difference between the two choice options.
5.1 Model Comparisons
In this section, we present the experimental results from our model comparisons. We first obtained embeddings from Arithmetic-GPT and LLaMA3 models, evaluating versions of the model that were pretrained on each of our four distinct synthetic datasets as well as a version without any training. Specifically, we extracted embeddings for the expected values of the two options, denoted as $e_{A}$ and $e_{B}$. Additionally, we obtained embeddings for the difference in expected values, denoted as $e_{A-B}$. All embeddings were derived from the representation in the final layer before the autoregressive prediction. We then performed a logistic regression using $e_{A}$, $e_{B}$, and $e_{A-B}$ as independent variables, with human choice probabilities as the dependent variable. Adjusted $R^{2}$ values were used for all logistic regressions, including those for Arithmetic-GPT, LLaMA3, and EV results. All other $R^{2}$ values were reported as the squared Pearson’s $r$ between model predictions and human data. These results indicate that the embeddings from the Arithmetic-GPT model pretrained on ecologically valid synthetic datasets most accurately capture human choices. This model also outperforms the embeddings obtained from the LLaMA3 model (the 7th row of Table 3), suggesting that pretraining on synthetic datasets is sufficient to create a strong correspondence between LLMs and human decision-making. However, the LLaMA3 model performs comparably to Arithmetic-GPT in predicting intertemporal choices. The log probabilities from the same LLaMA3 model perform poorly in comparison to human data (the 6th row of Table 3), replicating previous findings using choice rates reported from LLaMA1 (Binz & Schulz, 2023a). The $R^{2}$ values between Arithmetic-GPT pretrained on uniform and ecological synthetic datasets are small yet consistent. This may be due to the limited range tested in the human datasets.
To benchmark the performance of Arithmetic-GPT models in explaining human data, we directly fit behavioral models and MLPs on each of the human choice datasets. Prospect theory and the hyperbolic discounting model are leading behavioral models in risky and intertemporal choices, respectively (Kahneman & Tversky, 2013; Laibson, 1997). The behavioral models have interpretable mechanisms and contain few free parameters. However, they do not explain human data as well as the embeddings from both LLaMA3 and Arithmetic-GPT, although they still outperform a simple choice model based on EV difference. Moreover, prospect theory and the hyperbolic discounting model do not generalize to both risky and intertemporal choices, resulting in N/A values in their respective rows of Table 3.
In contrast, fitting an MLP directly on human data potentially reveals the ceiling performance of any model in explaining the data (Agrawal et al., 2020; Peterson et al., 2021). Except for the intertemporal choice tasks, MLPs outperform all other models. It is important to note that MLP training was based on task features rather than text descriptions that simulated participants’ experiences or arithmetic equations that mimic a rational agent’s computations. Consequently, these differences in input formats could also lead to diverging performance. The arithmetic format of intertemporal choice may provide a better fit for human data. Indeed, there is increasing evidence from experimental economics suggesting that the complexity of discounting values, even in the absence of actual payoff delays, influences intertemporal choices (Enke et al., 2023). We include a robustness check using 10-fold cross-validation in Appendix B.
5.2 Implicit functions of probabilities, values, and discount factors
@@
Figure 2: Embeddings from ecologically pretrained Arithmetic-GPT for inputs including (A) probabilities, (B) values, and (C) discount factors. Inputs are shown along the horizontal axes and embeddings are shown on the vertical axes. The embeddings, shown as black dots, were reduced to 1D using multidimensional scaling. Embeddings for probabilities and discount factors are normalized between 0 and 1 (see Appendix D for details). The red curves represent the best-fitting behavioral economic models: (A) the probability weighting function from PT with best-fitting $\gamma=0.58$, (B) the utility function from PT with best-fitting $\alpha=0.42,\beta=0.45,\lambda=1.4$, and (C) the hyperbolic discount function with best-fitting $k=0.08$.
We find that the embeddings from ecologically pretrained Arithmetic-GPT replicate classical findings from behavioral economics, including value and probability weighting functions from the prospect theory (Tversky & Kahneman, 1992) and the hyperbolic discounting function (Laibson, 1997). Specifically, probabilities close to 0.5 are more similar to each other than to probabilities close to either 0 or 1 (see Figure 2A and Equation 4). Embeddings for values illustrate concavity for positive values, convexity for negative values, and a steeper slope for negative values than for positive values (see Figure 2B and Equation 3). These features reflect risk aversion for gains, risk seeking for losses, and loss aversion (Kahneman & Tversky, 2013; Tversky & Kahneman, 1992). Moreover, embeddings for the discount factor demonstrate that distant times are more similar (see Figure 2C and Equation 6), enabling a present-bias (Meier & Sprenger, 2010). However, embeddings from Arithmetic-GPT pretrained on non-ecological datasets (i.e., ablated ecological, uniform, and ablated uniform) and those from LLaMA3 exhibit fewer human-like distortions in their implicit functions (see Appendix E for details).
For the probability weighting function, the curvature parameter ($\gamma$) of human participants was found to be 0.61 for gains and 0.69 for losses (Tversky & Kahneman, 1992). Follow-up replications by (Wu & Gonzalez, 1996) found $\gamma$ values around 0.7. Moreover, the utility curvature parameters ($\alpha$ and $\beta$) typically range between 0.5 and 0.9, while the loss aversion parameter ($\lambda$) is approximately 2.25 (Tversky & Kahneman, 1992). Regarding human intertemporal choices, typical $k$ values range between 0.01 and 0.1 for small magnitude studies (Odum, 2011). Comparing to these best-fitting parameters from human experiments, we observe that the implicit functions derived from the embeddings of Arithmetic-GPT also quantitatively match those observed in humans (see Figure 2). The implicit functions, however, exhibit discontinuities, suggesting that the smooth functions derived from human theories may not be directly applicable to explain the embeddings of Arithmetic-GPT.
@@ 6 Discussion
We introduced an approach to transforming language models into cognitive models by pretraining them on tasks that mirror the computational process performed by a rational agent. Specifically, we pretrained Arithmetic-GPT to calculate expected values using ecologically distributed probabilities and values. This pretraining allows the model to better capture human risky and intertemporal choices compared to classical behavioral models and, in some cases, even surpass the performance of the general-purpose LLaMA3 model. These results suggest that the observed cognitive biases in LLMs may stem from the training task, architecture, and numerical reasoning abilities, without requiring natural language supervision. These results have implications for a number of questions in cognitive science and machine learning, although we also note the limitations of our current work.
@@
Language Models as Cognitive Models. There is a growing research effort focused on exploring LLMs as scientific tools for understanding the human mind (Binz & Schulz, 2023a; Frank, 2023b; Horton, 2023). Despite the challenges outlined in Section 2, LLMs offer unique opportunities for investigating cognitive processes in ways that are not feasible with human participants. Recent studies have even demonstrated that fine-tuning a LLaMA1 model on human data allows the model to outperform classical models in explaining this data (Binz & Schulz, 2023a). While the LLaMA series’ model architecture and weights are publicly available, researchers lack access to the training data, which makes crucial scientific inference practically impossible. In contrast, we explicitly manipulate the training data for our Arithmetic-GPT, thereby uncovering a key factor that contributes to the model’s ability to explain human choices. This targeted approach, however, comes at the cost of the broader versatility inherent in off-the-shelf LLMs, which are designed to handle a wider range of tasks. Notably, language models, including Arithmetic-GPT, do not directly model human cognitive processes. Instead, our work suggests that these models are better suited for probing the computational problems that human cognition aims to solve.
@@
Bayesian Models of Cognition, Meta-learning, and Pre-training. Bayesian models of cognition have been instrumental in understanding human performance across a variety of cognitive tasks by providing optimal solutions to the inductive inference problem. Recently, it has been argued that Bayesian models and neural network models can be viewed as complementary to one another (Griffiths et al., 2023). Neural networks that are meta-trained have also been shown to exhibit properties consistent with Bayesian models (Lake & Baroni, 2023; McCoy & Griffiths, 2023). Moreover, pretraining a neural network model for improved performance on downstream tasks can be seen as a form of meta-learning or the acquisition of useful inductive biases, similar to Bayesian models (Hsu et al., 2018). Our work makes the implicit priors learned by Arithmetic-GPT more explicit by specifying the synthetic dataset on which it was pre-trained.
@@
Computationally Equivalent Tasks for Cognitive Modeling. Modeling human cognition is challenging because the hypothesis space of possible cognitive mechanisms that can explain human data equally well is vast. This makes principles of rationality desirable, as assuming people are rational in some sense greatly constrains the hypothesis space, thereby making scientific inference more effective. However, human rationality has been a subject of debate in economics and psychology for over a century (von Neumann & Morgenstern, 1944; Kahneman & Tversky, 2013; Simon, 1997). While significant progress has been made in understanding human rationality (e.g., Lieder & Griffiths, 2020; Gershman et al., 2015), the advent of LLMs seems to challenge the need for rational theories. Simply training LLMs to predict the next word appears sufficient to produce human-like behaviors, suggesting that we can model human cognition without the constraints imposed by rational theories. However, our experimental results suggest an alternative route to enhancing the correspondence between behaviors produced by LLMs and humans: pretraining on computationally equivalent tasks that a rational agent would need to master. Future research should investigate the impact of different assumptions about the nature of rationality on task content and distributions, and explore whether there are more effective assumptions for pre-training models to explain human behavior.
@@
Implications for Theories of Human Risk and Time Preferences. The success of Arithmetic-GPT in explaining human risky and intertemporal choices has significant implications for theoretical work on human risk and time preferences. While the Homo economicus portrayal of human beings as perfectly rational and self-interested agents has been inadequate in describing human choices (Gigerenzer & Gaissmaier, 2011; Kahneman & Tversky, 2013), existing behavioral models that deviate from rationality do not generalize well across different task domains. For instance, it is challenging to use prospect theory to model time preferences or to use the hyperbolic discounting model to explain risk preferences. In contrast, Arithmetic-GPT has demonstrated substantial transferability across task domains. The same model, in principle, can be adapted to other judgment and decision-making tasks such as social choice, probability judgment, and belief updating. The key factor enabling language models to generalize across a wide range of tasks is the presence of a language interface, which underpins a significant range of cognitive tasks.
@@
The Importance of Training Data Disclosure. Our results demonstrate that the training data used for LLMs is crucial for understanding their emergent capabilities and the inductive biases they acquire during pretraining. Adjusting the distribution of, or ablating, the training data can significantly affect the degree to which LLMs correspond with human behaviors. These findings suggest that existing off-the-shelf LLMs, whether proprietary or open-weight, including the GPT series (Brown et al., 2020) and the LLaMA series (Touvron et al., 2023), are less effective as models of human cognition. This is primarily because their training data is rarely disclosed, making it difficult for scientists to control for data contamination and thereby precisely identify the sources of human-like behaviors in these models.
@@
Limitations and Future Research. While we have made progress in addressing many challenges associated with using language models as cognitive models, some issues remain unresolved. To address the data gap between LLMs and human learners, we limited the scope of our synthetic dataset to the arithmetic of expected value calculations. Despite this, LLMs still require a substantial amount of training data to perform arithmetic accurately within a limited range of input values. Moreover, it is unrealistic for human learners to process 1M randomly generated mathematical equations, as Arithmetic-GPT did, to acquire the skill of computing expected values. Further research is needed to continue bridging this data gap (but see Appendix F for an initial attempt).
@@
Additional ablation studies could be performed on model architectures and training objectives. Our work is fundamentally limited to autoregressive training and a decoder-only transformer architecture. We believe that alternative training mechanisms and model architectures could potentially yield better embeddings from language models. One area for future research is estimating the lower bounds on model size necessary to achieve a certain level of correspondence with human behavior. Another area of future work involves leveraging interpretability techniques to distill novel cognitive mechanisms from Arithmetic-GPT.
@@ Appendix A Prompts
System: You are participating in a gambling game where you will be shown two options, Gamble A and Gamble B. Your task is to choose between the two. You must select one option.
User: Gamble A offers a 10% chance to win \$10 and a 90% chance to lose \$12. Gamble B offers a 40% chance to lose \$13 and a 60% chance to win \$22. Please limit your answer to either ‘A’ or ‘B’.
@@
System: You are participating in an intertemporal choice experiment. You will be presented with two options, A and B, and your task is to choose between them. You must select one option.
User: Option A offers \$80 after 30 days. Option B offers \$13 immediately. Please limit your answer to either ‘A’ or ‘B’.
@@ Appendix B Robustness Check
We supplement our main results with a robustness check using 10-fold cross-validation. Each dataset of human choices was randomly partitioned into a 90% training set and a 10% validation set. For the gershman20 dataset, a stratified train/validation split was employed due to the certainty equivalence design of the experiment (Gershman & Bhui, 2020), ensuring that choices made within the same problem were grouped together.
Embeddings from Arithmetic-GPT models and LLaMA-3-70B-Instruct were fitted using logistic regression with a LASSO penalty on the training set. MLPs were fitted to task features in the training set. Performance for all models was assessed on the validation set, with $R^{2}$ calculated as the squared Pearson’s $r$. This process was repeated 10 times, each with a random train/validation split. The mean and standard error (SE) of $R^{2}$ are summarized in Table A1, demonstrating a replication of our main results presented in Table 3. Finally, we conducted a model comparison using the entire choices13k dataset, including the decision-from-feedback trials (see Table A2). We observe that LLaMA3 can outperform MLP on the agrawal23 dataset. While the two neural network models use different task formats as inputs, which may contribute to the differences in performance, it is also possible that the dataset lacks sufficient statistical power to reliably distinguish between the two models.
Table A1: Cross-validation results. Propotion of the variance in human choices explained ($R^{2}$) on the validation set for each model. Numbers in parentheses represent standard errors from 10-fold cross-validation.
@@
Model Training data choices13k cpc18 gershman20 agrawal23
Arith.-GPT Synthetic (unif.) 56.4% (0.4%) 46.3% (3.1%) 56.7% (1.1%) 81.5% (0.4%)
@@
Note. The bold numbers represent the top-performing variants of the Arithmetic-GPT models, as measured by $R^{2}$. In cases where multiple models are highlighted, the differences in $R^{2}$ values are not statistically significant at the $p=0.05$ level.
Table A2: Proportion of the variance in human choices explained ($R^{2}$) on the entire choices13k dataset for each model. Bold numbers indicate the best models within each group.
Model Training data Entire choices13k
Arith.-GPT Synthetic (unif.) 64.0%
Arith.-GPT Synthetic (unif. abl.) 49.2%
Arith.-GPT Synthetic (eco.) 64.3%
Arith.-GPT Synthetic (eco. abl.) 52.1%
LLaMA3 (txt.) Undisclosed 14.6%
LLaMA3 (txt emb.) Undisclosed 65.6%
LLaMA3 (arith.) Undisclosed 60.6%
Prospect theory Human choices 45.6%
Expected value None 40.2%
MLP Human choices 73.8%
Appendix C Learning Curve
Figure A1: (A) Training loss decreases over the course of training epochs for Arithmetic-GPT trained on ecological synthetic dataset. (B) Histogram displaying the differences between the top-1 responses of the pretrained Arithmetic-GPT model and the actual expected values in the ecological synthetic dataset.
Appendix D Dimensionality Reduction Details
We used multidimensional scaling to reduce the dimensionality of embeddings to 1D, with Euclidean distance serving as the dissimilarity metric for forming the 1D manifolds. Embeddings of probabilities and discount factors were normalized using max-min normalization to ensure the resultant values fell within the range of $[0,1]$.
Appendix E Additional Implicit Functions
Figure A2: Visualizations of embeddings from LLaMA-3-70B-Instruct.
Figure A3: Visualizations of embeddings from Arithmetic-GPTs pretrained on different synthetic datasets. (A) Uniform synthetic dataset. (B) Ablated uniform synthetic dataset. (C) Ablated ecological synthetic dataset.
Appendix F Variations in Model Sizes, Data Quantity, and Data Distributions
In this section, we explore how adjustments to key hyperparameters of Arithmetic-GPT impact the quality of its pretrained embeddings. Specifically, we investigate the effects of reducing the model’s hidden size from 320 (as reported in the main text) to 104 and 16. Moreover, we examine the influence of decreasing the size of the ecological synthetic dataset, reducing it from 1M mathematical equations (as reported in the main text) to subsets of 100K and 10K equations. Apart from these modifications, we adhered to the same pretraining procedure and evaluated the resulting embeddings on the two largest datasets, choices13k for risky choices (Peterson et al., 2021) and agrawal23 (Agrawal et al., 2023) for intertemporal choices, using 10-fold cross-validation.
The results, summarized in Table A3, reveal an intriguing pattern: reducing the size of the ecological synthetic dataset to as few as 10K equations does not significantly impair the ability of the 10M model’s embeddings to predict human choices. It is important to note that Arithmetic-GPT models were never exposed to human data during pretraining. However, significant reductions in the model size – particularly 30K parameters – lead to a decline in performance, highlighting the critical role of model capacity in embedding quality.
Table A3: Variants of Arithmetic-GPT models pretrained on ecological synthetic datasets of varying sizes. In each cell, $R^{2}$ results (in percentage) are reported as choices13k (left of ‘/’) and agrawal23 (right of ‘/’). Numbers in parentheses indicate standard errors obtained from 10-fold cross-validation.
Model Sizes
10M (hidden size=320) 1M (hidden size=104) 30K (hidden size=16)
1M 57.7 (0.4) / 83.7 (0.3) 54.7 (0.6) / 83.4 (0.3) 30.6 (0.4) / 56.8 (0.4)
Data quantity 100K 57.2 (0.6) / 83.6 (0.4) 55.5 (0.5) / 81.6 (0.4) 31.4 (0.8) / 63.4 (0.2)
10K 57.7 (0.5) / 83.6 (0.3) 50.0 (0.5) / 84.1 (0.3) 29.1 (0.6) / 63.0 (0.4)
Note. The bold numbers represent the top-performing variants of the Arithmetic-GPT models, as measured by $R^{2}$. In cases where multiple models are highlighted, the differences in $R^{2}$ values are not statistically significant at the $p=0.05$ level.
We conducted an exploratory analysis using the original Arithmetic-GPT model architecture and synthetic datasets of 1M equations to investigate the effects of varying data distributions. Specifically, we independently manipulated the distributions of probabilities and values. For probability distributions, we considered Beta(0.27, 0.27) (ecological), Beta(1, 1) (uniform), and Beta(2, 2). For value distributions, we examined power-law distributions with exponents of 0 (uniform), -0.945 (ecological), and -2. All Arithmetic-GPT models were randomly initialized and trained following the procedure outlined in Section 3.1.
As shown in Table A4, aligning probability distributions with ecological patterns, such as Beta(0.27, 0.27), generally enhances the quality of embeddings for predicting risky choices. Conversely, adopting a power-law distribution for values with an exponent of -2 tends to improve the model’s performance in predicting intertemporal choices.
Table A4: Variants of Arithmetic-GPT models pretrained on different data distributions. In each cell, $R^{2}$ results (in percentage) are reported as choices13k (left of ‘/’) and agrawal23 (right of ‘/’). Numbers in parentheses indicate standard errors obtained from 10-fold cross-validation.
Distribution of probabilities
Beta(0.27,0.27) Beta(1,1) Beta(2,2)
0 56.5 (0.5) / 82.5 (0.4) 56.4 (0.4) / 81.5 (0.4) 56.0 (0.3) / 77.5 (0.4)
Distribution of values (exponent of power-law) -0.945 57.7 (0.4) / 83.7 (0.3) 56.7 (0.3) / 78.4 (0.5) 56.5 (0.3) / 77.6 (0.4)
-2 55.7 (0.3) / 85.9 (0.3) 56.3 (0.5) / 82.6 (0.4) 55.5 (0.4) / 82.2 (0.4)
Note. The bold numbers represent the top-performing variants of the Arithmetic-GPT models, as measured by $R^{2}$.
Appendix G Computational Abilities of Language Models
We evaluate the abilities of LLaMA-3-70B-Instruct and Arithmetic-GPT in computing expected values using two newly generated test datasets. Both test sets consist of 20K synthetically generated equations involving expected value calculations. In one test set, the probabilities and values in the equations are sampled from a uniform distribution, while the other test set features probabilities and values derived from ecological distributions, as illustrated in Figure 1. The train-test relationships for these datasets are summarized in Table A5.
[…]
To evaluate LLaMA3’s ability to compute expected values, we directly prompted the model to solve arithmetic equations while setting the temperature to 0. The integer part of the expected values was extracted from LLaMA3’s responses by filtering out irrelevant tokens. The correct probability was calculated as the relative frequency of correctly predicting the integer part of the true expected values across 20K questions. As shown in Table A5, LLaMA3 performed poorly compared to Arithmetic-GPT, achieving only 20.97% and 38.79% accuracy on the uniform and ecological test sets, respectively.
The model’s behavioral performance on the ecological test set broadly predicts the ranking of its embeddings in modeling human data. However, directly interpreting a language model’s performance on arithmetic tasks in relation to model-fitting results that leverage its embeddings can be challenging. Therefore, further investigation is necessary to establish a robust connection between a language model’s arithmetic capabilities and the predictive power of its embeddings in predicting human choices.
Table A5: Evaluation of language models’ probabilities of generating correct expected values across different train-test distributions. Numbers in parentheses represent standard errors.
Model Training data Test data Correct probability
Arith.-GPT Synthetic (unif.) Uniform 0.8261 (0.0023)
Ecological 0.9336 (0.0014)
Arith.-GPT Synthetic (eco.) Uniform 0.9388 (0.0013)
Ecological 0.9780 (0.0007)
LLaMA3 Undisclosed Uniform 0.2097
Ecological 0.3879
Appendix H Implementation Details